# NestedFP: High-Performance, Memory-Efficient Dual-Precision Floating Point Support for LLMs

### v2

@@ 5.4 Impact on SLO Attainment
Methodology. In this section, we present a case study demonstrating how NestedFP can improve SLO attainment. We implement a simple load-aware mode switching policy for NestedFP that, in each iteration, selects FP8 mode when the number of tokens to process exceeds a predefined threshold (1024 in this experiment); otherwise, it selects FP16 mode. The evaluation is conducted using Llama 3.1–70B on a cluster equipped with four NVIDIA H100 GPUs. As input, we use 1,000 requests sampled from an Azure LLM inference trace while varying the request rate. We measure p90 TTFT, p90 TPOT, and SLO attainment under two criteria: loose and tight. The loose SLO is defined as $\leq 10\times$ the average single-request latency, and the tight SLO is defined as $\leq 5\times$ the average single-request latency, following the methodology of a previous study sola_llm .
@@ Appendix A Evaluation Details
We conducted our evaluation using the Mistral Nemo Base 2407 and Mistral Small 24B Base 2501 models. For accuracy assessment, we employed the Math Verify Metric on the Minerva Math task, the Open LLM Leaderboard configuration (5-shot, multiple choice) for the MMLU Pro task, and a zero-shot evaluation setting for the BBH task. Kernel performance was measured using NVIDIA Nsight Compute. In order to get optimal kernel, we padded activation in the multiple of tile dimension $T_{m}$. During performance measurement, we flushed all caches using the cache-control option, but did not fix the GPU clock frequency using the clock-control option.
Appendix B Extended Kernel Performance for NestedFP
Figure 8: Performance comparison between our CUTLASS baseline and NestedFP kernel across 14 unique $(N,K)$ GEMM shapes. Each subplot represents a different shape configuration, with batch size $(M)$ on the $x$-axis and latency on the $y$-axis.
Figure 8 compares our method against the CUTLASS FP16 baseline across all 14 GEMM shapes for the four evaluated models: Llama 3.1 8B, Mistral Nemo, Phi-4, and Mistral Small. The matrix dimension $M$ is varied in increments of 32, ranging from 32 to 2048. Our results show that NestedFP consistently incurs only moderate overhead across all GEMM configurations, with an average performance difference of 6.1%. The per-graph average overhead ranges from 4.3% to 7.2%, demonstrating stable performance regardless of matrix shape. This consistency highlights the effectiveness of our kernel-level optimizations.
Appendix C Extended End-to-End Throughput Evaluation for NestedFP
@@
Figure 9: Extended end-to-end throughput evaluation across different input/output token configurations. Each subplot shows throughput comparison for FP16 and NestedFP across four models (Llama 3.1 8B, Mistral Nemo, Phi-4, Mistral Small) under different request size settings: (a) 32 input, 512 output tokens (b) 1024 input, 512 output tokens (c) 32 input, 32 output tokens and (d) 1024 input, 32 output tokens.
We extend our end-to-end evaluation to further demonstrate the robustness of NestedFP. Figure 9 presents results for four additional input/output configurations: (32, 512), (1024, 512), (32, 32), and (1024, 32). Across all settings, NestedFP exhibits minimal overhead compared to the FP16 baseline, with an average degradation of 5.52% and a maximum of 7.69%. On a per-model basis, the average overhead is 5.97% for LLaMA 3.1 (8B), 6.67% for Mistral-Nemo (12B), 5.51% for Phi-4 (14B), and 5.05% for Mistral-Small (24B).
Our evaluation setup mirrors the configuration used in the main throughput experiments. Throughput was computed as the total number of generated tokens divided by the end-to-end latency for each test. All tests were conducted on a single NVIDIA H100 GPU using the vLLM 0.8.5 V1 engine with PyTorch 2.7.0+cu12.6 for both the baseline and our proposed NestedFP implementations. Chunked prefill, enabled by default in the vLLM V1 engine, was used throughout the evaluation. This feature segments the prefill phase into smaller chunks, which are processed concurrently with decoding tokens within the same batch. At each iteration, the GEMM operation shape $(M,N,K)$ is determined based on the number of tokens grouped in that step, with $M$ denoting the total token count. We set the maximum number of batched tokens to 8192 for the throughput test.
@@ Appendix D CUTLASS FP16 GEMM Kernel
D.1 CUTLASS FP16 Baseline and PyTorch Kernel Comparison
Figure 10: Performance comparison between PyTorch kernel and our CUTLASS baseline across 14 unique $(N,K)$ GEMM shapes. Each subplot represents a different shape configuration, with batch size $(M)$ on the $x$-axis and latency on the $y$-axis.
To validate the robustness of our CUTLASS baseline, we compare its performance against the cuBLAS kernels employed by PyTorch 2.6.0 with CUDA 12.4. Using a weak baseline could underestimate the overhead introduced by NestedFP, leading to overly optimistic conclusions about its practical efficiency. Figure 10 provides a detailed comparison across 14 $(N,K)$ configurations, extracted from the linear layers of the four models evaluated in our main experiments—Llama 3.1 8B, Mistral Nemo, Phi-4, and Mistral Small. For each configuration, we sweep the batch size $M$ from 32 to 2048 in increments of 32, covering the typical range encountered in LLM inference workloads.
Our results show that the FP16 baseline closely matches the performance of the cuBLAS implementation, with an average performance difference of just 1.8% across all 14 GEMM shapes. In 11 of the 14 configurations, our baseline achieves higher average performance than PyTorch, with performance differences reaching up to 3.3%. For the remaining three shapes—corresponding to smaller GEMMs—PyTorch slightly outperforms our implementation, with differences ranging from 0.3% to 0.9%, well within the bounds of normal GPU performance variability. These results confirm that our CUTLASS baseline is not only competitive but often exceeds production-level kernel performance, ensuring a reliable and fair comparison when evaluating NestedFP’s overhead.
@@ D.2 Kernel Search Space
The kernel design space consists of two variants: non-cooperative and cooperative kernels. For non-cooperative kernels, we perform a grid search over tile dimensions with $T_{m}\in{16,32,64,128,256}$, $T_{n}\in{64,128,256}$, and $T_{k}\in{64,128,256}$. Cooperative kernels adopt larger tiles with $T_{n}\in{128,256}$, and configurations that fail to compile are excluded. Non-cooperative kernels use a cluster shape of $(1,1,1)$, whereas cooperative kernels use $(2,1,1)$. We employ the persistent scheduler for non-cooperative kernels, and both the persistent and Stream-K [25] schedulers for cooperative kernels.
Appendix E Broad Applicability of NestedFP
Model GEMM1 GEMM2 GEMM3 GEMM4 Total
CodeLlama 7B 96/96 32/32 64/64 31/32 223/224 (99.6%)
CodeLlama 13B 120/120 40/40 80/80 37/40 277/280 (98.9%)
Gemma 3 4B 207/264 64/88 123/176 34/34 429/563 (76.2%)
Gemma 3 12B 249/306 78/102 151/204 48/48 527/661 (79.7%)
Gemma 3 27B 291/348 92/116 179/232 62/62 625/759 (82.3%)
Llama 3.1 8B 96/96 32/32 64/64 32/32 224/224 (100.0%)
Llama 3.1 70B 224/240 80/80 141/160 78/80 523/560 (93.4%)
Mistral Nemo 12B 120/120 40/40 80/80 40/40 280/280 (100.0%)
Mistral Small 24B 120/120 40/40 80/80 40/40 280/280 (100.0%)
Phi-3.5 Mini 26/32 31/32 31/32 24/32 112/128 (87.5%)
Phi-4 14B 40/40 38/40 40/40 28/40 146/160 (91.2%)
Qwen 3 8B 108/108 35/36 72/72 34/36 249/252 (98.8%)
Qwen 3 14B 120/120 40/40 80/80 38/40 278/280 (99.3%)
Qwen 3 32B 192/192 63/64 127/128 56/64 438/448 (97.8%)
Table 4: Layer-wise applicability of NestedFP across models. Format X/Y indicates X applicable layers out of Y total layers for each GEMM type.
Our scheme is broadly applicable to a wide range of LLMs beyond those evaluated in the main experiments. In Table 4, we analyze ten additional models that exhibit the key characteristic: the majority of their weight values have absolute magnitudes less than or equal to 1.75. For each model, we examine four distinct GEMM shapes corresponding to different linear layer types: GEMM1 (QKV projections), GEMM2 (output projections), GEMM3 (MLP gate/up projections), and GEMM4 (MLP down projections). Each entry is shown as X/Y, indicating that X out of Y layers of that type satisfy the weight magnitude constraint (absolute value $\leq$ 1.75).
Our analysis indicates that most models exhibit high applicability rates. The Gemma 3 family shows the lowest rates (76.2%–82.3%), primarily due to multi-modal projection layers containing weights with significantly higher magnitudes (up to 26.25). Among the various GEMM types, GEMM4 (MLP down projections) tends to include more layers exceeding the threshold in models such as Phi-4 14B (70.0% applicability) and Qwen 3 32B (87.5%). While the majority of models have maximum weight values below 3.0, notable outliers include Llama 3.1 70B (maximum value of 93.0) and the Gemma 3 series (maximum value of 26.25). These extreme values are typically confined to a small number of layers and have limited impact on the overall applicability of our method. The consistently high applicability rates across diverse model families affirm the robustness and generalizability of our approach’s core assumption regarding weight distributions.

### v3

@@ 5.4 Impact on SLO Attainment
Methodology. In this section, we present a case study demonstrating how NestedFP can improve SLO attainment. We implement a simple load-aware mode switching policy for NestedFP that, in each iteration, selects FP8 mode when the number of tokens to process exceeds a predefined threshold (1024 in this experiment); otherwise, it selects FP16 mode. The evaluation is conducted using Llama 3.1–70B on a cluster equipped with four NVIDIA H100 GPUs. As input, we use 1,000 requests sampled from an Azure LLM inference trace while varying the request rate. We measure p90 TTFT, p90 TPOT, and SLO attainment under two criteria: loose and tight. The loose SLO is defined as $\leq 10\times$ the average single-request latency, and the tight SLO is defined as $\leq 5\times$ the average single-request latency, following the methodology of a previous study Hong et al. (2025).
@@ Appendix A Evaluation Details
Experiments were conducted on NVIDIA H100 GPU with vLLM 0.8.5 (V1 engine) and PyTorch 2.7.0+cu12.6. For accuracy assessment, we use the Math Verify metric on Minerva Math, the Open LLM Leaderboard configuration for MMLU Pro, and a zero-shot setting for BBH. The $M$ dimension is padded to multiples of $T_{m}$ only for kernel microbenchmarks and is disabled in the vLLM integration.
Appendix B Extended Kernel Evaluation
Figure 8: Comparison of kernel performance between the CUTLASS baseline and NestedFP.
Figure 8 compares NestedFP with the CUTLASS FP16 baseline across 14 $(N,K)$ GEMM shapes for four models: Llama 3.1 8B, Mistral Nemo, Phi-4, and Mistral Small. We vary $M$ from 32 to 2048 in increments of 32. Overall, NestedFP incurs a moderate overhead, with an average of 6.38% relative to the FP16 baseline.
Appendix C Extended End-to-End Throughput Evaluation
@@
Figure 9: FP16 inference throughput comparison across four models under different input/output token configurations.
Figure 9 reports end-to-end throughput for (input, output) token pairs of (32, 512), (1024, 512), (32, 32), and (1024, 32). Across all settings, NestedFP achieves throughput comparable to the FP16 baseline. Chunked prefill was enabled, and the maximum number of batched tokens was set to 8192.
@@ Appendix D CUTLASS FP16 GEMM Kernel
D.1 CUTLASS Baseline and PyTorch Kernel Comparison
Figure 10: Comparison of kernel performance between CUTLASS baseline and PyTorch.
Figure 10 compares kernel performance between the CUTLASS FP16 baseline and PyTorch 2.6.0 (CUDA 12.4). We evaluate 14 $(N,K)$ GEMM shapes from four models: Llama 3.1 8B, Mistral Nemo, Phi-4, and Mistral Small, sweeping $M$ from 32 to 2048 in increments of 32. Overall, the CUTLASS baseline achieves performance comparable to PyTorch.
@@ D.2 Kernel Search Space
We build upon CUTLASS 3.6.0. In CUTLASS, four warps are grouped into a warp group. We evaluate two variants: a non-cooperative kernel (one warp group for TMA and one for MMA) and a cooperative kernel (one warp group for TMA and two for MMA). For non-cooperative kernels, we grid-search tile dimensions with $T_{m}\in\{16,32,64,128,256\}$, $T_{n}\in\{64,128,256\}$, and $T_{k}\in\{64,128,256\}$. For cooperative kernels, we vary only $T_{n}$, setting $T_{n}\in\{128,256\}$. We exclude configurations that fail to compile. Non-cooperative kernels use a cluster shape of $(1,1,1)$, whereas cooperative kernels use $(2,1,1)$. We use the persistent scheduler for non-cooperative kernels, and consider both the persistent and Stream-K [25] schedulers for cooperative kernels.
Appendix E Extended Applicability Analysis
Model # Layers # Applicable Ratio (%)
CodeLlama 7B 224 223 99.6
CodeLlama 13B 280 277 98.9
Gemma 3 4B 563 429 76.2
Gemma 3 12B 661 527 79.7
Gemma 3 27B 759 625 82.3
Llama 3.1 8B 224 224 100.0
Llama 3.1 70B 560 523 93.4
Mistral Nemo 12B 280 280 100.0
Mistral Small 24B 280 280 100.0
Phi-3.5 Mini 128 112 87.5
Phi-4 14B 160 146 91.2
Qwen 3 8B 252 249 98.8
Qwen 3 14B 280 278 99.3
Qwen 3 32B 448 438 97.8
Table 4: Layer-wise applicability of NestedFP across models.
Table 4 presents an applicability analysis across a broader range of models. We report the total number of layers, the number of applicable layers satisfying the weight-magnitude constraint ($|w|\leq 1.75$), and the corresponding ratio. Overall, most models exhibit high applicability across layers.